Two liters of an ideal gas is collected at 25 °C and 745 mmhg. What is the volume at standard sea level conditions?
1
Expert's answer
2020-11-10T10:00:42-0500

At standard sea level conditions the temperature is 15oC, and the pressure is 760 mmHg.

25oC = 298.15 K

15oC = 288.15 K

According to the combined gas law,

"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}."

Therefore,

"V_2=\frac{P_1V_1T_2}{T_1P_2}=\frac{745mmHg\times2.00L\times288.15K}{298.15K\times760mmHg}=1.89L"


Answer: 1.89 L